which is the greatest 3 digit number exactly divisible by 8,10,12
Answer by Mathtut(3670)

960....which is 8 times 10 times 12. lets test numbers of above that
:
since it has to be divisible by 12 lets add 12 to 960
:
960+12=972...10 doesnt work here
:
972+12=984...10 doesnt work here
:
984+12=996...10 doesnt work here
:
any other additions of 12 would take us to 4 digits.